National Defence Academy (NDA) · Topic Collection

Quantitative Aptitude PYQs

Arithmetic, number systems, percentages, ratios, profit and loss, time and work, algebra, geometry, mensuration, and trigonometry.

2520 Total Questions21 Official PapersYears: 2015–2025Search in Quantitative Aptitude
Showing questions 341–360 of 2520Page 18 of 100
NDA-II 2024 (Maths) Official Paper (Held On: 01 Sept, 2024) (Previous Year Paper)·Mathematics
View full paper

Question 341

National Defence Academy (NDA) 2024• Mathematics• Mathematics

Let x - 3y + 4 = 0 and 2x - 7y + 8 = 0 be two lines of regression computed from some bivariate data. If byx and bxy are regression coefficients of lines of regression of y on x and x on y respectively, then what is the value of bxy + 7byx?

  1. AOption (a)

    -2

  2. BOption (b)

    1

  3. COption (c)

    2

  4. DOption (d)

    5

View answer & explanation

Correct answer

Answer: (d)D.

5

Why this answer is correct

Explanation:

Lines of regression of x on y

⇒x - 3y + 4 = 0

⇒x = 3y – 4

⇒ bxy = 3

Line of regression of y on x

⇒ 2x – 7y+8 = 0

⇒ y =\(\frac{2}{7}x + \frac{8}{7}\)

⇒ byx = 2/7

Now

⇒ b xy + 7b yx  =  3 + 7× 2 /7 = 5

∴ Option (d) is correct

NDA-II 2024 (Maths) Official Paper (Held On: 01 Sept, 2024) (Previous Year Paper)·Mathematics
View full paper

Question 342

National Defence Academy (NDA) 2024• Mathematics• Mathematics

The mean of n observations

1, 4, 9, 16 ...., n2 is 130. What is the value of n?

  1. AOption (a)

    18

  2. BOption (b)

    19

  3. COption (c)

    20

  4. DOption (d)

    21

View answer & explanation

Correct answer

Answer: (b)B.

19

Why this answer is correct

Explanation:

Mean  = \(\frac{1+ 4+ 9 ... n^2}{n}\)

⇒ 130 = \(\frac{(n(n+1)(2n+1)}{6n}\)

⇒ 780 = 2n2 + 3n + 1

⇒ 2n2 + 3n – 779 =0

⇒ 2n2 + 41n – 38n – 779 = 0

⇒ n(2n + 41) – 19(2n + 41) = 0 

⇒ (2n + 41) (n – 19) = 0

⇒ x = 19 ..... ( n = \(-\frac{41}{2}\) is not possible)

∴ Option (b) is correct.

NDA-II 2024 (Maths) Official Paper (Held On: 01 Sept, 2024) (Previous Year Paper)·Mathematics
View full paper

Question 343

National Defence Academy (NDA) 2024• Mathematics• Mathematics

Three distinct natural numbers chosen at random from 1 to 10. What is the probability that they are consecutive?

  1. AOption (a)

    1/12

  2. BOption (b)

    3/40

  3. COption (c)

    1/15

  4. DOption (d)

    7/120

View answer & explanation

Correct answer

Answer: (c)C.

1/15

Why this answer is correct

Explanation:

⇒n(S) = 10C3

Favourable outcomes = (1, 2, 3), (2, 3, 4), (3, 4, 5), (4, 5, 6), (5, 6, 7), (6, 7, 8), (7, 8, 9), (8, 9, 10)

⇒n(E) = 8

⇒ P(E) = \(\frac{n(E)}{n(S)} = \frac{8}{^{10}C_3}\)

= \(\frac{8\times3\times2\times1}{10\times9 \times8} =\frac{1}{15}\)

∴ Option (c) is correct

NDA-II 2024 (Maths) Official Paper (Held On: 01 Sept, 2024) (Previous Year Paper)·Mathematics
View full paper

Question 344

National Defence Academy (NDA) 2024• Mathematics• Mathematics

A, B, C are three mutually exclusive and exhaustive events associated with a random experiment. If 3P(B) = 4P(A) and 3P(C) = 2P(B), then what is P(A) equal to?

  1. AOption (a)

    7/29

  2. BOption (b)

    8/29

  3. COption (c)

    9/29

  4. DOption (d)

    10/29

View answer & explanation

Correct answer

Answer: (c)C.

9/29

Why this answer is correct

Explanation:

A, B and C are mutually exclusive and exhaustive events:

⇒  P(A) + P(B) + P(C)= 1.....(i)

Also 

⇒ 3P(B) = 4P(A)

⇒ P(A) = \(\frac{3}{4} P(B)\)

and 3P(C) = 2P(B) 

⇒ P(C) = \(\frac{2}{3}P(B)\)

From (i) 

⇒ \(\frac{3}{4}P(B) +P(B)+ \frac{2}{3}P(B) =1\)

⇒ \(\frac{9P(B)+12P(B)+8P(B)}{12} = 1 \)

⇒ \(\frac{29P(B)}{12} =1\)

⇒ P(A) = \(\frac{3}{4} \times \frac{12}{29} = \frac{9}{29}\)

∴ Option (c) is correct.

NDA-II 2024 (Maths) Official Paper (Held On: 01 Sept, 2024) (Previous Year Paper)·Mathematics
View full paper

Question 345

National Defence Academy (NDA) 2024• Mathematics• Mathematics

A die has two faces with number 4, three faces with number 5 and one face with number 6. If the die is rolled once, then what is the probability of getting 4 or 5?

  1. AOption (a)

    1/3

  2. BOption (b)

    2/3

  3. COption (c)

    5/6

  4. DOption (d)

    1/2

View answer & explanation

Correct answer

Answer: (c)C.

5/6

Why this answer is correct

Explanation:

Number on faces on die are 4, 4, 5, 5, 5 and 6.

⇒n(S) = 6

P (getting 4 or 5) = 5/6

∴ Option (c) is correct

NDA-II 2024 (Maths) Official Paper (Held On: 01 Sept, 2024) (Previous Year Paper)·Mathematics
View full paper

Question 346

National Defence Academy (NDA) 2024• Mathematics• Mathematics

A box contains 2 black, 4 yellow and 6 white balls. Three balls are drawn in succession with replacement. What is the probability that all three are of the same colour?

  1. AOption (a)

    1/6

  2. BOption (b)

    1/36

  3. COption (c)

    1/12

  4. DOption (d)

    5/12

View answer & explanation

Correct answer

Answer: (a)A.

1/6

Why this answer is correct

Explanation:

Black = 2, yellow = 4 and white = 6

⇒ Total Balls = = 2 + 4 + 6 = 12

P (all three are of the same colour).

= \(\frac{2^3}{12^3}+ \frac{4^3}{12^3} +\frac{6^3}{12^3} = \frac{1}{6}\)

∴ Option (a) is correct

NDA-II 2024 (Maths) Official Paper (Held On: 01 Sept, 2024) (Previous Year Paper)·Mathematics
View full paper

Question 347

National Defence Academy (NDA) 2024• Mathematics• Mathematics

A can hit a target 5 times in 6 shots, B can hit 4 times in 5 shots and C can hit 3 times in 4 shots. What is the probability that A and C may hit but B may lose?

  1. AOption (a)

    1/8

  2. BOption (b)

    1/6

  3. COption (c)

    1/4

  4. DOption (d)

    1/3

View answer & explanation

Correct answer

Answer: (a)A.

1/8

Why this answer is correct

Explanation:

Here, P(A) = 5/6 , P(B) = 4/5 , P(C) = 3/4

Also

P(B) = \(1 - \frac{4}{5} = \frac{1}{5}\)

Now

⇒ \(P ( A\cap \overline{B} \cap C) = P(A).P(B).P(C)\)

= \(\frac{5}{6}\times \frac{1}{5}\times \frac{3}{4} = \frac{1}{8}\)

∴ Option (a) is correct.

NDA-II 2024 (Maths) Official Paper (Held On: 01 Sept, 2024) (Previous Year Paper)·Mathematics
View full paper

Question 348

National Defence Academy (NDA) 2024• Mathematics• Mathematics

The letters of the word ZOOLOGY are arranged in all possible ways. What is the probability that the consonants and vowels occur alternatively?

  1. AOption (a)

    6/35

  2. BOption (b)

    3/35

  3. COption (c)

    2/35

  4. DOption (d)

    1/35

View answer & explanation

Correct answer

Answer: (d)D.

1/35

Why this answer is correct

Explanation:

The word given is ZOLOGY. We need to find the probability that the consonants and vowels occur alternatively in all possible arrangements.

  • Vowels: O, O, O
  • Consonants: Z, L, G, Y

There are 3 vowels and 4 consonants.

The total number of ways the letters of the word "ZOOLOGY" can be arranged (without any restrictions) is:

Formula: Total arrangements = n! / r1! r2! ... rk!

Where:

  • n = Total number of letters = 7
  • r1 = Frequency of O = 3

So, the total number of arrangements is:

Total arrangements = 7! / 3! = 5040 / 6 = 840

We need to place the consonants and vowels alternately:

  • Consonants will occupy the odd positions (1, 3, 5, 7).
  • Vowels will occupy the even positions (2, 4, 6).

The number of ways to arrange the consonants (Z, L, G, Y) in the 4 odd positions is:

4! = 24

The number of ways to arrange the vowels (O, O, O) in the 3 even positions is:

3! / 3! = 1

Thus, the number of favorable arrangements is:

Favorable arrangements = 4! × (3! / 3!) = 24 × 1 = 24

The probability is given by the ratio of favorable arrangements to the total arrangements:

Probability = Favorable arrangements / Total arrangements = 24 / 840 = 1 / 35

The probability that the consonants and vowels occur alternatively is 1 / 35.

NDA-II 2024 (Maths) Official Paper (Held On: 01 Sept, 2024) (Previous Year Paper)·Mathematics
View full paper

Question 349

National Defence Academy (NDA) 2024• Mathematics• Mathematics

A natural number x is chosen at random from the first 100 natural numbers. What is the probability that x2 + x > 50?

  1. AOption (a)

    93/100

  2. BOption (b)

    47/50

  3. COption (c)

    24/25

  4. DOption (d)

    23/25

View answer & explanation

Correct answer

Answer: (b)B.

47/50

Why this answer is correct

Explanation:

⇒n(S) = 100 

Numbers from 7 to 100 satisfy x2 + x > 50

Then, n(E)= 100 – 6 = 94

⇒ \(P(E) = \frac{n(E)}{n(S)} = \frac{94}{100} = \frac{47}{50}\)

∴ Option (b) is correct

NDA-II 2024 (Maths) Official Paper (Held On: 01 Sept, 2024) (Previous Year Paper)·Mathematics
View full paper

Question 350

National Defence Academy (NDA) 2024• Mathematics• Mathematics

What is the mean deviation of the first 10 natural numbers?

  1. AOption (a)

    2

  2. BOption (b)

    2.5

  3. COption (c)

    3

  4. DOption (d)

    3.5

View answer & explanation

Correct answer

Answer: (b)B.

2.5

Why this answer is correct

Explanation:

The mean of 10 natural numbers 

⇒ \(\overline{x} = \frac{1+2+3+4+....10}{10}\) 

= \(\frac{10\times11}{2\times10}\) = 5.5

Mean deviation = \(\frac{[1- 5.5] [2-5.5] [3- 5.5] ... [10- 5.5]}{10} = 2.5\)

∴ Option (b) is correct.

NDA-II 2024 (Maths) Official Paper (Held On: 01 Sept, 2024) (Previous Year Paper)·Mathematics
View full paper

Question 351

National Defence Academy (NDA) 2024• Mathematics• Mathematics

Let \(\rm \Sigma_{i=1}^9x_i^2=855\) If M is the mean and σ is the standard deviation of  x1, x2, x3.....x9 then what is the value of M2 + σ2?

  1. AOption (a)

    100

  2. BOption (b)

    95

  3. COption (c)

    90

  4. DOption (d)

    85

View answer & explanation

Correct answer

Answer: (b)B.

95

Why this answer is correct

Explanation:

Given:

\(\rm \Sigma_{i=1}^9x_i^2=885\)

Variance σ = \(\rm \Sigma_{i=1}^9x_i^2 - (mean)^2\)

= \(\frac{855}{9} -(M)^2\)

\(x = {-b \pm \sqrt{b^2-4ac} \over 2a}\)

⇒ σ + M2 = \(\frac{855}{9} = 95\)

∴ Option (b) is correct.

NDA-II 2024 (Maths) Official Paper (Held On: 01 Sept, 2024) (Previous Year Paper)·Mathematics
View full paper

Question 352

National Defence Academy (NDA) 2024• Mathematics• Mathematics

The mean of the series x1, x2...xn is x̅ If xn is replaced by k, then what is the new mean?

  1. AOption (a)

    x̅ - xn + k

  2. BOption (b)

    \(\rm \frac{n\bar x-\bar x+k}{n}\)

  3. COption (c)

    \(\rm \frac{\bar x-x_n-k}{n}\)

  4. DOption (d)

    \(\rm \frac{n\bar x-x_n+k}{n}\)

View answer & explanation

Correct answer

Answer: (d)D.

\(\rm \frac{n\bar x-x_n+k}{n}\)

Why this answer is correct

Explanation:

Given:

\(\overline x = \frac{x_1+x_2+x_3+......... x_n}{n}\)

⇒ \(x_1+x_2+X_3+......x_n= n\overline x \)

⇒ Mean when xn is replaced by k

⇒ New mean = \(\frac{x_1+x_2+x_3+......... x_{n-1}+k}{n}\)

= \(\frac{n\overline x -x_n+k}{n}\)

∴ Option (d) is correct

NDA-II 2024 (Maths) Official Paper (Held On: 01 Sept, 2024) (Previous Year Paper)·Mathematics
View full paper

Question 353

National Defence Academy (NDA) 2024• Mathematics• Mathematics

A fair coin is tossed till two heads occur in succession. What is the probability that the number of tosses required is less than 6?  

  1. AOption (a)

    5/64

  2. BOption (b)

    15/32

  3. COption (c)

    31/64

  4. DOption (d)

    19/32

View answer & explanation

Correct answer

Answer: (d)D.

19/32

Why this answer is correct

Explanation:

A fair coin is tossed till two heads occur in succession.

We are asked to find:

P(the number of tosses required is less than 6)

This means we need to compute the probability that the first occurrence of two consecutive heads (HH) happens at or before the 5th toss.

Valid sequences where HH first appears at the end are:

HH (length 2)

THH (length 3)

TTHH, HTHH (length 4)

TTTHH, THTHH, HTTHH (length 5)

So the total probability is:

P = P(HH) + P(THH) + P(TTHH) + P(HTHH) + P(TTTHH) + P(THTHH) + P(HTTHH)

= \((\frac{1}{2})^2 + (\frac{1}{2})^3 + 2 \times (\frac{1}{2})^4 + 3 \times (\frac{1}{2})^5\)

= \(\frac{1}{4} + \frac{1}{8} + \frac{2}{16} + \frac{3}{32}\)

= \(\frac{8 + 4 + 4 + 3}{32} = \frac{19}{32}\)

∴ Option (4) is correct.

NDA-II 2024 (Maths) Official Paper (Held On: 01 Sept, 2024) (Previous Year Paper)·Mathematics
View full paper

Question 354

National Defence Academy (NDA) 2024• Mathematics• Mathematics

Urn A contains 2 white and 2 black balls while urn B contains 3 white and 2 black balls. One ball is transferred from urn A to urn B and then a ball is drawn out of urn B. What is the probability that the ball is white?

  1. AOption (a)

    11/20

  2. BOption (b)

    7/12

  3. COption (c)

    3/5

  4. DOption (d)

    1

View answer & explanation

Correct answer

Answer: (b)B.

7/12

Why this answer is correct

Explanation:

Given:

Urn A contains 2 white and 2 black balls while urn B contains 3 white and 2 black balls. 

⇒ E1 = One white ball transferred from A to B; then urn A contains white = 1 and b = 2 and urn B contains white = 4 and black = 2

⇒ E2 = One black ball transferred from A to B. Then urn A contains white = 2 and black = 1 and urn B contains white = 3 and black = 3

⇒ F = One white ball is drawn from urn B

⇒ \(P(F) =P(E_1). P(\frac{F}{E_1})+ P(E_2).P(\frac{F}{E_2})\)

= \(\frac{2}{4}\times\frac{4}{6}+\frac{2}{4}\times\frac{3}{6} = \frac{7}{12}\)

∴ Option (b) is correct.

NDA-II 2024 (Maths) Official Paper (Held On: 01 Sept, 2024) (Previous Year Paper)·Mathematics
View full paper

Question 355

National Defence Academy (NDA) 2024• Mathematics• Mathematics

For two events A and  B, P(A) = P(A|B) = 0.25 and P(BIA) = 0.5. Which of the following are correct?

I. A and B are independent.

II. P(Ac ∪ Bc) = 0.875

III. P(Ac ∩ Bc) = 0.375

Select the answer using the code given below.

  1. AOption (a)

    I and II only

  2. BOption (b)

    II and III only

  3. COption (c)

    I and III only

  4. DOption (d)

    I, II and III

View answer & explanation

Correct answer

Answer: (d)D.

I, II and III

Why this answer is correct

Explanation:

Given:

\(P(A) = P(\frac{A}{B}) = 0.25\)

and \(P(\frac{B}{A}) = 0.5\)

I. \(P(\frac{B}{A}) = \frac{P(A∩ B)}{P(A)}\)

⇒ P(A∩B) = P(A) P(B|A)

⇒ P(A∩B) = 0.25 × 0.5 = 0.125

Now

⇒ \(P(\frac{B}{A}) = \frac{P(A∩ B)}{P(B)}\)

⇒ \(P(B)= \frac{P(A∩ B)}{P(\frac{A}{B})}\)

⇒ \(P(B) = \frac{0.125}{0.25} = 0.5\)

Now, P(A).P(B) = 0.25 × 0.5 = 0.125 = P(A∩B)

Thus  A and B are independent

II. \(P(\overline A\cup \overline B ) = 1 – P(A ∩ B)\)

= 1 – 0.125 = 0.875

III. \(P(\overline A∩ \overline B ) = 1 – P(A \cup B)\)

= 1 – [P(A) + P(B) – P(A ∩ B)

=  1 – [0.25 + 0.5 – 0.125]

= = 1 – 0.625 = 0.375

So all statements I, II, and III are correct.

∴ Option (d) is correct.

NDA-II 2024 (Maths) Official Paper (Held On: 01 Sept, 2024) (Previous Year Paper)·Mathematics
View full paper

Question 356

National Defence Academy (NDA) 2024• Mathematics• Mathematics

Two perfect dice are thrown. What C is the probability that the sum of the numbers on the faces is neither 9 nor 10?

  1. AOption (a)

    1/36

  2. BOption (b)

    5/36

  3. COption (c)

    7/36

  4. DOption (d)

    29/36

View answer & explanation

Correct answer

Answer: (d)D.

29/36

Why this answer is correct

Explanation:

⇒n(S) = 6 × 6 = 36

⇒Sum 9 = (3, 6), (4, 5), (5, 4), (6, 3)

⇒ Sum 10 = (4, 6), (5, 5), (6, 4)

⇒P (sum 9 or 10) = 7/36

Then, P (sum neither 9 nor 10) = \(1- \frac{7}{36} =\frac{29}{36}\)

∴ Option (d) is correct

NDA-II 2024 (Maths) Official Paper (Held On: 01 Sept, 2024) (Previous Year Paper)·Mathematics
View full paper

Question 357

National Defence Academy (NDA) 2024• Mathematics• Mathematics

The occurrence of a disease in an industry is such that the workers have 20% chance of suffering from it. What is the probability that out of 6 workers chosen at random, 4 or more will suffer from the disease?

  1. AOption (a)

    53/3125

  2. BOption (b)

    63/3125

  3. COption (c)

    73/3125

  4. DOption (d)

    83/3125

View answer & explanation

Correct answer

Answer: (a)A.

53/3125

Why this answer is correct

Explanation:

Given

n = 6

⇒ p = P (suffering from disease) = 20/100 = 1/5

⇒ q = P (not suffering from disease) = 1 - 1/5 = 4/5

⇒ p(x ≥ 4) = p(x = 4) + p(x = 5) + p(x = 6)

⇒ \(^6C_4(\frac{1}{5})^4 (\frac{4}{5})^2 +^6C_5 (\frac{1}{5})^5(\frac{4}{5})^1+^6C_6(\frac{1}{5})^6\)

= \(\frac{6!}{4!2!}\times\frac{16}{5^6}+ \frac{6!}{5!}\times\frac{4}{5^6}+ \frac{1}{5^6}\) 

= \(\frac{265}{15625} = \frac{53}{3125}\)

∴ Option (a) is correct

NDA-II 2024 (Maths) Official Paper (Held On: 01 Sept, 2024) (Previous Year Paper)·Mathematics
View full paper

Question 358

National Defence Academy (NDA) 2024• Mathematics• Mathematics

Three perfect dice are rolled. Under the condition that no two show the same face, what is the probability that one of the faces shown is an ace (one)?

  1. AOption (a)

    5/9

  2. BOption (b)

    2/3

  3. COption (c)

    1/3

  4. DOption (d)

    1/2

View answer & explanation

Correct answer

Answer: (d)D.

1/2

Why this answer is correct

Explanation:

⇒n(S) = 6 × 5 × 4

⇒ n(E) = n (one of the faces shown is an ace.)

=  3(1 × 5 × 4) = 3 × 5 × 4 

⇒ \(P(E) = \frac{3\times 5\times 4}{6\times 5\times4} = \frac{1}{2}\)

∴ Option (d) is correct

NDA-II 2024 (Maths) Official Paper (Held On: 01 Sept, 2024) (Previous Year Paper)·Mathematics
View full paper

Question 359

National Defence Academy (NDA) 2024• Mathematics• Mathematics

Three perfect dice D1, D2 and D3 and are rolled. Let x, represent the numbers on D1, D2 and D3 respectively. What is the number of possible outcomes such that x < y < z?

  1. AOption (a)

    20

  2. BOption (b)

    18

  3. COption (c)

    14

  4. DOption (d)

    10

View answer & explanation

Correct answer

Answer: (a)A.

20

Why this answer is correct

Explanation:

Possible outcomes x < y < z

= = (1, 2, 3) (1, 2, 4), (1, 2, 5), (1, 2, 6) (1, 3, 4) (1, 3, 5) (1, 3, 6), (1, 4, 5), (1, 4, 6), (1, 5, 6), (2, 3, 4),(2, 3, 5) (2, 3, 6), (2, 4, 5), (2, 4, 6), (2, 5, 6), (3, 4, 5), (3, 4, 6) (3, 5, 6), (4, 5, 6)

Thus, Total possible outcomes = 20

∴ Option (a) is correct.

NDA-II 2024 (Maths) Official Paper (Held On: 01 Sept, 2024) (Previous Year Paper)·Mathematics
View full paper

Question 360

National Defence Academy (NDA) 2024• Mathematics• Mathematics

In a binomial distribution, if the mean is 6 and the standard deviation is √2, then what are the values of the parameters n and p respectively?

  1. AOption (a)

    18 and 1/3

  2. BOption (b)

    9 and 1/3

  3. COption (c)

    18 and 2/3

  4. DOption (d)

    9 and 2/3

View answer & explanation

Correct answer

Answer: (d)D.

9 and 2/3

Why this answer is correct

Explanation:

Given:

⇒ Mean x = np = 6....(i)

⇒ SD = \(\sqrt{npq} = \sqrt2\)...(ii)

Dividing (ii) by (i), we get

⇒ \(\frac{npq}{np} = \frac{2}{6}\)

⇒ q= 1/3

Now

⇒ p = 1 – q = 1- 1\3 = 2/3

From (i)

⇒\(n\times \frac{2}{3} = 6\)

⇒n =9

∴ Option (d) is correct.

How should I prepare for Quantitative Aptitude in National Defence Academy (NDA)?

Solving previous year questions by topic helps identify recurring concepts, question patterns, and frequently asked sub-topics. Attempt each question, review the step-by-step explanation, and note down key formulas or facts for quick revision before the exam.